Fred wants to prepare a snack mix comprised of peanuts and dried fruit. Peanuts cost $5 a pound and dried fruit costs $7 a pound. Fred wants to spend exactly $45 and prepare 8 pounds of his snack mix. How many pounds of peanuts should fred purchase, and how many pounds of dried fruit should Fred purchase?
Answer by checkley77(12844)

7X+5(8-X)=45
7X+40-5X=45
2X=45-40
2X=5
X=5/2
X=2.5 POUNDS OF DRIED FRUIT IS NEEDED.
8-2.5=5.5 POUNDS OF PEANUTS IS NEEDED.
PROOF:
7*2.5+5*5.5=45
17.5+27.5=45
45=45